Find the length of the arc of the cal x square is equal to eat from the vertex to an activity of a letter rectum

1
Expert's answer
2022-01-11T13:46:37-0500

Find length of the arc of the parabola "x^2=4ay" measured from the vertex to one extremity of the Latus-Rectum

Latus-Rectum is a chord that passes through a focus and is parallel to the directrix


Let A be the vertex and L an extremity of the Latus-Rectum so that at A,x=0 and at L,x=2a.

Then:

"y=x^2\/(4a),y'=x\/(2a)"

focus: "(-a,0)"

directrix: "x=a"


arc "AL=\int^{2a}_0\sqrt{1+(y')^2}dx=\int^{2a}_0\sqrt{1+(\frac{x}{2a})^2}dx="


"x=2asinhu\to dx=2acoshudu"


"=\int 2acoshu\sqrt{4a^2sinh^2u+4a^2}du=4a^2\int cosh^2udu="


"=\frac{coshusinhu+u}{2}=(\frac{x\sqrt{(x/(2a))^2+1}}{2}+sinh^{-1}\frac{x}{2a})|^{2a}_0=a(\sqrt 2+sinh^{-1}1)"